3x-5y=-9
5x-6y=-8

Answer by Alan3354(69443)

3x-5y=-9
5x-6y=-8
Multiply both eqns to eliminate x
15x-25y=-45
15x-18y=-24
Subtract 2 from 1
0x -7y = -21
y = 3
Sub into eqn 1
3x -5*3 = -9
3x - 15 = -9
3x = 6
x = 2